I am having a really hard time answering this word problem. Here it is:
Two sisters washed the family car in 24 min. When they each washed the car alone, the older sister took 10 min longer to do the job than the younger sister. How would the younger sister take to wash the car alone?

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

Let the younger sister's time=x
the older sister's time is x+10
in 1 minute, the younger sister does 1/x of the car and the older sister does 1/(x+10) of the car.
in 24 minutes, they do the whole car together.
Therefore 24/x+24/(x+10)=1 (the whole car)
Multiply by x(x+10) both sides to clear fractions
24(x+10)+24x=x^2+10x, which is x(x+10)
x^2+10x=24x+240+24x
x^2-38x-240=0
x=(1/2)(38+/-sqrt(38^2+960); sqrt term is of (2404)=49.03
Therefore, the one positive root is (1/2)87.03=43.52 min
43.5 min for the younger sister ANSWER
53.5 min for the older sister
(24/43.5)=0.55
(24/53.5)=0.45

Answer by ikleyn(52786)

.
For more similar solved joint work problems see the lesson
    - Using quadratic equations to solve word problems on joint work
in this site.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this textbook under the topic "Rate of work and joint work problems" of the section "Word problems".